# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v2

@@ 5. Recruited Participants
We received over 3000 responses from 48 countries. See Appendix G for sample characteristics. Out of a total of five trials, participants finished four trials on average (Table 2), suggesting the experiment was engaging. Most participants came from the Creative Mornings newsletter or r/InternetIsBeautiful (Table 3 for source counts and categorization). The sample was 50% women, 43% men, 4% non-binary, 3% not disclosed, 1% self-described. The mean age was 34.92 (SD = 10.86). Regarding AI, the sample was 48% neutral, 28% excited, 24% concerned. Participants said they were more creative than 57.86% (SD = 26.66) of AI and 58.67% (SD = 23.65) of humans. See Appendix Figure 11 for kernel density plots. Users from neutral interest groups who were concerned about AI tended to have low self-reported creativity.
6. Outcome Measures
We have three outcome measures (idea diversity, creativity, and AI adoption) and three levels of analysis (local, evolution, and global). See Table LABEL:big_table. The local level measures outcomes at the level of an individual trial (e.g., how a submitted response relates to example responses). The evolution level measures the rate of change of outcome variables with respect to the trial number in the response chain (i.e., experiment iteration). The global level compares all submitted responses in a condition to each other. For all pairwise comparisons, we use a Holm-Bonferroni adjustment for multiple comparisons. For idea diversity and AI adoption, we scale the dependent variable (cosine distance or cosine similarity, respectively) by 100 for easier interpretation.
Table 4. We measure three outcome measures (idea diversity, creativity, AI adoption) and three levels of analysis (local, global, and evolution). If a level of analysis is not appropriate for an outcome measure, we put a ‘Not applicable’ in that cell. All ideas are embedded using SBERT.
Local Global Evolution
Creativity How creative is the submitted response? This is measured by the prediction of the classifier from Organisciak et al. (2022). Not applicable Does the creativity of submitted ideas change over time? This is measured by the slope of the response chain’s trial number (i.e., iteration in the response chain) on creativity (the metric from Organisciak et al. (2022)).
Idea Diversity How different is a participant’s response from example responses? This is measured by the median pairwise semantic distance between ideas a participant sees and their response. How diverse were all the participant’s ideas in a condition? This is measured by the median pairwise distance between all submitted ideas in a condition. Do ideas become more different from each other as the experiment goes on? We first measure the median pairwise distance (‘idea diversity’) of ideas at each trial number (i.e., iteration in the response chain). We then measure the slope of the trial number on idea diversity.
AI Adoption How similar is a participant’s response to AI example responses? This is measured by the maximum semantic distance between a participant’s response and AI examples. Not applicable Not applicable
6.1. Local Level
@@
Outcomes at the local level—the level of an individual trial—are useful for two reasons. First, this level shows how a participant’s response relates to the examples they see. Second, this level lets us model whether individual differences moderate the effect of either disclosure or transparency. For each of our local outcomes, we have a baseline model that uses crossed random intercepts to account for the multilevel structure of the experiment. The first random intercept is for participants, accounting for clustering due to repeated measures. This random intercept is then crossed with a second random intercept for response chains, which we nest inside of {[condition], item} combinations. [^10] In R syntax, the random effect structure was ... + (1|ParticipantID) + (1|ItemCondition/ResponseChainID); See Figure 4 for a visual explanation of how response chains are nested in items and conditions. Models were fit in the lme4 R package. We computed profile likelihood confidence intervals for coefficients using the confint R package. We used estimated marginal means (emmeans R package) to conduct model-adjusted F-tests, linear contrasts, predictions, and pairwise comparisons. We apply Holm-Bonferroni adjustments to pairwise comparison p-values. Our baseline ‘local’ model is:
$\text{variable}_{ijk}$ $=\beta_{0}+\beta_{1}\text{condition}_{j}+\beta_{2}\text{CreativityHuman}_{i}+\beta_{3}\text{AiRelCreate}_{i}+$
$\phantom{=}\beta_{4}\text{AiFeeling}_{i}+\beta_{5}\text{InterestGroup}_{i}+\beta_{6}\text{ConditionOrder}_{ijk}$
$\phantom{=}\beta_{7}\text{LogDuration}_{ijk}+\beta_{8}\text{nSeedsPresent}_{ijk}+\beta_{9}\text{TrialNo}_{jk}$
$\phantom{=}u_{0i}+v_{0jk}+e_{ijk}$
@@
  • AiRelCreate is constructed as (self-perceived creativity to humans) - (self-perceived creativity to AI). Note that this is an implicit measure of AI’s creativity relative to humans. For example, if you say you are more creative than 40% of humans and 60% of AI, then AiRelCreate = -20, as the implicit belief is AI is less creative (-20 percentile points) than humans. Conversely, if you say you are more creative than 20% of AI but 50% of humans then the implicit belief is humans are more creative (50% - 30% = +20).
@@
We balanced interest in testing experimental hypotheses that conditions differed by subgroups with caution around an over-fitted model. We considered interactions between the treatment condition and four potential moderators: self-perceived human creativity ($CreateHuman$), AI - Human creativity ($AiRelCreate$), feeling towards AI ($AIFeeling)$, and interest group ($InterestGroup$). We first conducted likelihood ratio tests to test if adding each moderator improved our baseline model. Moderators were kept only if they significantly improved the fit $(p<0.05)$. See Appendix Table 11 for retained moderators. Then, we used emmeans to probe and interpret moderating effects.
6.2. Global

### v3

@@ 5.1. Recruited Participants
Figure 7. Median pairwise distance of submitted ideas in a condition from Monte Carlo analysis. There was more global diversity of ideas in both high AI exposure conditions than in the control condition.
Figure 8. High exposure to AI ideas increased the rate of change in idea diversity. ‘Idea diversity’ was measured as the cosine distance between submitted ideas.
5.2. High AI exposure increases collective diversity
Intuitively, global idea diversity measures how different submitted ideas are in each condition. We measured global idea diversity by looking at the median pairwise distance between neural embeddings of the pool of ideas that were submitted by participants in each condition (Appendix B.2). As a non-parametric measure of effect size, we calculate Cliff’s Delta ($\delta$) of these pairwise differences, which ranges from -1 to 1. A value of 0 indicates no difference between the two conditions, +1 indicates values from the first condition are always larger, and -1 indicates the opposite.
We find (Figure 7) that both the [High Exposure, Disclosed] ($\delta=0.31;\text{p}=0.001)$ and [High Exposure, Undisclosed] $(\delta=0.26;\text{p}=0.001)$ conditions had more global idea diversity than the control condition. But of the low exposure conditions, only the [Low Exposure, Undisclosed] $(\delta=0.11;\text{p}=0.001)$ condition had higher global diversity than the control condition, with a much smaller effect size than the high exposure conditions. P-values were computed with permutation tests incorporating a conservative adjustment (Ojala and Garriga, 2009) (adding 1 to numerator and denominator) and a Holm-Bonferroni correction for multiple comparisons. As robustness checks, we ran this analysis with alternative measures of idea diversity (SI 8) and the results are directionally similar. We report results for the median pairwise distance as it provides the most conservative estimate.
5.3. High AI exposure increases the speed at which idea diversity develops
We measured the evolution of diversity by looking at the diversity (Appendix B.3) of submitted ideas at each iteration in the experiment and seeing if—in some conditions—diversity developed at a faster rate. We find that high AI exposure increases not only the average levels (Figure 7) of collective idea diversity but also the rate of change in idea diversity (Figure 8). As with global idea diversity, different metrics yielded similar regression coefficients (SI 8.3). In the control condition, idea diversity decreased over trials ($\beta=-0.39$, $t(349)=-2.23$, 95% CI = [0.73,0.05][-0.73,-0.05][ - 0.73 , - 0.05 ], $p=0.03$). That is, submitted ideas were becoming more similar to each other as the experiment went on. Relative to the control condition, however, the slope of idea diversity with respect to trial number was more positive for the [High Exposure, Undisclosed] condition ($\beta=0.53$, $t(349)=2.2$, 95% CI = [0.06,0.99][0.06,0.99][ 0.06 , 0.99 ], $p=0.03$) and the [High Exposure, Disclosed] condition ($\beta=0.57$, $t(349)=2.37$, 95% CI = [0.1,1.03][0.1,1.03][ 0.1 , 1.03 ], $p=0.02$). The rate of change in idea diversity for the low AI exposure conditions did not differ from the rate of change in the control condition. Thus, we conclude that high exposure to AI ideas increased the rate of idea diversity relative to the no-AI, control condition.
5.4. Seeing AI ideas does not affect creativity
Conditions did not differ by creativity of ideas, $F(4,19.86)=0.12,p=0.97$. We used the emmeans [^5] https://cran.r-project.org/web/packages/emmeans/index.html joint-tests function to assess main effects of condition after a mixed model (Appendix B.1 for modeling details; SI 9 for regression results). Similarly, no individual treatment condition coefficient differed from zero where the baseline was the control condition. We also tested whether the evolution of creativity differed by condition. We conducted a likelihood ratio test on whether adding a [trial number (as an integer) X condition] interaction would improve the fit of a model. The likelihood ratio test indicated that allowing for this temporal interaction did not significantly improve the model fit $(\chi^{2}(4)=6.52,p=0.16)$—suggesting conditions did not differ on average or with respect to the temporal evolution of creativity. All results are based on mixed models (Appendix B.1) with crossed and nested intercepts accounting for non-independence in the experiment.
5.5. How AI Adoption differs depending on task and self-perceived creativity
As a secondary outcome to creativity and diversity, we measured whether AI adoption (the extent to which people submit similar ideas to AI examples) differs depending on the task and person. Prior work found that people may rely more on AI for more difficult tasks (Roemmele, 2021; Goddard et al., 2014). Other work posited that attitudes towards a technology are a major factor when studying human-AI collaboration. (Gero et al., 2022). We measured three attitudinal variables: Attitude toward AI, self-perceived creativity, and perceived creativity of AI. For our individual-level model of adoption, we only included variables as moderators if they significantly improved the fit of a no-moderator, baseline model (Appendix B.1). We operationalized AI adoption as the maximum similarity (following Roemmele (2021)) between a submitted idea and AI example ideas a participant saw. The logic for using the maximum is that insofar as a person is influenced by an idea, it would likely be a single one (Roemmele, 2021).
Figure 9. Higher creativity participants adopted AI ideas based on exposure and were not influenced by AI disclosure (i.e., whether AI ideas were labeled as such). ‘AI Adoption’ was measured by the similarity of a submitted idea to AI example ideas.
Figure 10. When AI ideas were disclosed as such, participants were more likely to adopt AI ideas for more difficult prompts.
5.5.1. People who identify as creative are less influenced by AI disclosure
[…]
5.5.2. Participants may adopt AI ideas more for harder prompts
We find that when AI ideas are disclosed, participants are more likely to adopt the ideas of AI for difficult AUT prompts (Figure 10). The rank-order correlation between difficulty and adoption was $\rho=0.8$ for the [High Exposure, Disclosed] condition but only $\rho=0.3$ for the [High Exposure, Undisclosed] condition. That is, when people were told ideas were from AI, they were more likely to adopt AI ideas if the prompt was difficult. But since we employed only five items, this finding is speculative. We encourage future work to follow up on this finding.
6. Discussion
Against the backdrop of widespread LLM exposure, we asked: How does exposure to ideas generated by LLMs affect the creativity, diversity, and evolution of human ideas? To answer this, we conducted a large-scale experiment where participants submitted ideas in response to the Alternate Uses Task (a measure of creativity where people brainstorm novel uses of an item) after seeing a set of example ideas. The human examples were from prior participants in the same experimental condition, and the AI ideas were from ChatGPT. The evolving aspect of our experiment, that ideas in a condition feed forward to subsequent trials in that condition, captures the fact that ideas build upon prior ideas. This lets us approximate the cultural evolution effects of AI. Here are three takeaways from our experiment.
6.1. AI makes ideas different but not better
On the other hand, the presence of AI ideas increased the diversity of human ideas. This is consistent with work suggesting collaborating with AI leads to more diverse or unexpected outputs (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019) and inconsistent with other work that finds collaborating with LLMs decreases diversity (Anderson et al., 2024; Doshi and Hauser, 2024; Dell’Acqua et al., 2023). But we highlight that our study is testing passive exposure to AI ideas and not active engagement with AI ideas. Our setup—passive exposure to AI ideas, scattered amongst human ones—maps onto how many users now experience AI ideas. Hence, it may be that active engagement with LLMs decreases content diversity but simply seeing these ideas as “sparks” (Gero et al., 2022) increases content diversity. And because many more people may be passively exposed to LLM outputs than actively engaging with LLMs, the effect of passive exposure is important to understand.
Another key difference between our study and many similar studies is the brainstorming context. Prior studies typically examine individual participants working directly with AI systems. Our study simulates group brainstorming where participants see ideas from prior group members. In group contexts, AI ideas may serve as additional “voices” that prevent convergence in human-only groups. This suggests AI’s effect on ideas may depend on whether it’s used for individual ideation versus collective brainstorming. But since ideas we see affect the ideas we create (Nijstad and Stroebe, 2006), there is an element of collective (in a broad sense) brainstorming even to individual brainstorming.
Crucially, high AI exposure increased both average amounts of diversity and rates of change in idea diversity. The latter result is especially important. Small differences in rates of change can accumulate to create large aggregate differences over time. This points to a dynamic that may be missed by static experimental designs. Future work—both simulations and dynamic experiments—can explore the implications of this increase in collective idea diversity unaccompanied by an increase in average individual creativity. For instance, can this dynamic generate “innovation”?
Our finding around the evolution of diversity (Figure 8) is instructive and points to a likely mechanism: Seeing other people’s ideas reduced idea diversity in the control condition over time, as successive participants were “converging” on a particular idea sequence. But then injecting AI ideas into the example set increased the diversity of submitted responses by partially “resetting” this convergence. This suggests off-the-shelf LLMs may be viewed as extensive, accessible idea banks that can counter groupthink in collaborative settings. Future research should examine the dynamics and boundary conditions of this potential mechanism’s effectiveness.
6.2. High creativity people are less influenced by the source label of ideas
For participants who viewed themselves as highly creative, adoption was driven by exposure and not disclosure (i.e., knowing ideas were from AI). But for lower-creativity participants, knowing the source of an idea did affect the adoption of that idea. Perhaps people high in self-reported creativity relied less on source cues when adopting ideas because they were more confident in their ability to judge an idea’s creative merit. Regardless, our results suggest that (self-reported) creative people will adopt ideas on the basis of their content. Knowing the source does not matter. In a world where humans have difficulty distinguishing if the content was AI-generated (Jakesch et al., 2022), these findings suggest people high in (self-reported) creativity may not be ‘duped’ into adopting AI ideas.
6.3. Participants may adopt AI ideas when the task is difficult
When AI ideas were labeled, participants were more likely to adopt AI ideas for difficult prompts rather than easy prompts. This finding is similar to what Roemmele (2021) observed, where seeing AI examples only influenced creative output when the task was difficult. Both findings are consistent with a theoretical account of task difficulty being associated with increased reliance on automation (Goddard et al., 2014). However, we emphasize that this finding should be taken as speculative (due to the small number of items) and not conclusive. We encourage future work to test whether users adopt AI ideas for more challenging creative tasks.
If it is the case that users turn to AI for difficult rather than trivial tasks, this would have several implications. While AI can potentially augment humans when they are stuck, researchers raised concerns over ‘model collapse’ (Shumailov et al., 2023)—the deteriorating performance of LLMs when trained on their outputs. If reliance on AI for creative tasks becomes routine, this may contribute to model collapse.
6.4. Limitations And Future Work
Our study has limitations that should be noted and can inform future work. First, we measured the effect of AI ideas for a single task. We chose this task because it is one of the most common creativity tasks (Abraham, 2016). But future work could explore other kinds of tasks—especially more complex ones.
Second, we had to operationalize ‘AI ideas’ in some concrete way (among a large array of choices), and future work can use a similar design but operationalize ‘AI ideas’ in a different way. We chose a commonly-used LLM at the time of the study, and our prompt was driven by ecological validity and prior work: We used a zero-shot prompt because that is what users would likely use, and the specific prompt we used was derived from prior research. We chose not to vary prompts because we didn’t want to increase the complexity of an already complex experiment. Future work could explore whether different AI elicitation procedures yield different results. This could involve different models, prompts, LLM-based ideation systems (e.g.,  (Rick et al., 2023; Pu et al., 2025)), or “simulated social ensembles” (Ashkinaze et al., 2025). A specific design choice we made is that LLMs were not shown any example stimuli, but humans were. We did this to (A) reflect real-life LLM usage (which is often zero-shot, not few-shot) and (B) conceptually represent LLMs as a ‘bank’ of ideas that can break up convergence during group brainstorming. However, this specific operationalization leaves open the question of how LLMs versus humans may come up with different ideas if they were given the same example stimuli.
Third, our finding about AI adoption and task difficulty is based on only five AUT items (tire, pants, shoe, table, bottle) and should be considered speculative, not definitive. We encourage future work to explore this hypothesis further. Fourth, future work should test if alternative classifiers or ways of conceiving variables yield different results. For idea diversity and AI adoption, we addressed this problem by showing that conceptually similar ways of measuring variables yielded qualitatively similar results. For our creativity measure, we used a highly accurate classifier (correlation with human judgments greater than $r=0.88$ for AUT items we used) trained for this exact task, for these exact AUT items (Appendix A.1). But of course, all models have some error, and future work based on this model propagates these errors. Incidentally, human judges of creativity only correlate with other human judges at $r=0.88$ (Organisciak et al., 2023), suggesting the classifier we used may be approaching ‘the approximate ceiling at which we could expect a model to correlate with human judgments’ (Organisciak et al., 2023, pg. 11) of creativity. Fifth, the classifier measures one facet of creativity: originality. Future work can also explore whether AI ideas have different effects on other facets of creativity such as fluency (how many ideas) or elaboration (how detailed an idea is). Sixth, we employed a non-representative sample of technology-interested users and creative professionals. While these two groups were most relevant to the phenomenon in question, our sample also limits generalizability. Seventh, our design captures some aspects of human-AI cultural evolution but not others; it is a stylized model that does not capture all the complexity of how AI can shape collective behavior (Farrell et al., 2025; Brinkmann et al., 2023). For example, future work could explore more complex transmission scenarios.
Despite these limitations, our work offers a large-scale, dynamic account of how seeing LLM ideas affects collective thought. This work opens the door to more dynamic experiments exploring the compounding effects of “AI in the culture loop”.
6.5. Conclusion: Passive exposure to AI ideas affects collective thought
Our experiment suggests that passive exposure to AI ideas—the kind of passive exposure we are now inundated with in a post-ChatGPT era—does affect collective thought. Even small effects are meaningful. This exposure is both pervasive and growing. But the effects of AI ideas are nuanced. Seeing AI ideas did not increase individual creativity, though it did increase collective diversity. The effects of AI ideas vary across individuals and tasks. There is still much to learn. We hope our study inspires more research on how passive exposure to AI ideas affects collective thought via the use of multiple-worlds experiments.
Acknowledgements.
We thank several of our colleagues for feedback, including: Joyce Chai, Harman Kaur, and Misha Teplitskiy. Apart from where described in the manuscript, we used generative AI for coding completions and light writing edits.
Appendix A Stimuli Construction
A.1. Choosing AUT Items
We had to choose a selection of items that people would brainstorm creative uses for. We chose five items for which the creativity classifier that we used had the highest accuracy. Previously, Organisciak et al. (2022) fine-tuned GPT-3 Davinci to predict the creativity of AUT items. We use a dataset from Organisciak et al. (2022) containing 20,121 responses from 2,025 participants, across 21 distinct AUT items and nine distinct studies (Organisciak et al., 2022). [^6] We obtained an early version of this dataset by direct correspondence with Dr. Organisciak on February 23, 2023; the code that Dr. Organisciak used to generate this dataset is available at https://github.com/massivetexts/llm_aut_study/blob/main/notebooks/Process_AUT_GT.ipynb Each response was graded for creativity by humans and normalized to a scale of 1-5. Then (Organisciak et al., 2022) fine-tuned GPT-3 Davinci to predict creativity scores. Here, fine-tuning involves providing {Input (an AUT response), Output (human rating)} pairs to a pre-trained LLM. Then the LLM adjusts its parameters to produce a similar output given an input, proxying human judgements. Overall, the fine-tuned GPT-3 classifier had a correlation [^7] We obtained scores for this classifier by downloading the zip file from (https://github.com/massivetexts/llm_aut_study/blob/main/results/evaluation.zip), then navigating to gt_main2/gpt-ft-davinci-1.csv of $r=0.81$ with human judgment. Accuracy varied by item (SI 1). For our experiment, we picked the five items for which the classifier had the highest accuracy: tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is “in-domain” for the classifier since we ask participants to do the same task for the same items the classifier was trained on.
A.2. Generating GPT Ideas
Prompt Construction
[…]
Word Length Prompt Experiment
We conducted a Monte Carlo experiment to confirm our modified zero-shot prompt resulted in responses with a similar word length to humans. Using parameters from Stevenson et al. (2022)’s experiments: For $n=1000$ trials, we fixed presence penalty and frequency penalty at 1, randomly chose a temperature (higher values lead to more randomness) in [0.65, 0.7, 0.75, 0.8], and randomly chose one of our 5 AUT items. For each trial, ChatGPT generated five ideas. The modified prompt resulted in responses with an average word length (M=4.44, SD = 1.34) much closer to human responses (M=4.56, SD = 4.97) than the original zero-shot prompt (M = 25.38, SD = 8.55). A permutation test (more details below) further shows that this difference was significant at $p<0.001$. We used the ideas generated by our modified zero-shot prompt as stimuli for the main experiment.
We conducted a permutation test of whether the limited-length zero-shot prompt (‘ZeroShotLimit’) yielded more similar word counts to the human responses than the original zero-shot prompt (‘ZeroShot’) from (Stevenson et al., 2022). Note that permutation tests are non-parametric, so they are robust to (e.g.) violations of normality. The observed difference is computed as:
$T^{obs}=$
$|Avg(\text{ZeroShot})-Avg(\text{HumanResponses})|-$
$|Avg(\text{ZeroShotLimit})-Avg(\text{HumanResponses})|$
where $Avg$ is the average word count of a response and the human responses are from (Organisciak et al., 2022).
Under the null hypothesis, we assume prompts do not differ in respective word counts, so we can swap the labels of ‘ZeroShot’ and ‘ZeroShotLimit’. To simulate draws from this null distribution, we permuted the labels of the source column 10000 times, each time calculating a new value for the difference of differences defined in $T$. The p-value is the proportion of permutations where the null $T$ is greater than or equal to the observed difference, $T^{obs}$. As in (Ojala and Garriga, 2009), we apply a conservative adjustment and add 111 to the numerator and denominator (this means we never get a p-value of 0).
$\text{p}=\frac{[\sum_{i=1}^{n}I(T^{null}>=T^{obs})]+1}{n+1}$
This p-value indicates the likelihood of observing a difference as extreme as our observed difference if the null hypothesis—that prompts did not differ in word counts—is true. We conclude the observed difference is significant.
Figure 11. To generate AUT ideas, we used the zero-shot prompt from (Stevenson et al., 2022) with an additional instruction at the end to match the mean length of human responses from prior work.
Appendix B Outcome Measures
B.1. Local Level
@@
Outcomes at the local level—the level of an individual trial—are useful for two reasons. First, this level shows how a participant’s response relates to the examples they see. Second, this level lets us model whether individual differences moderate the effect of either disclosure or exposure. For each of our local outcomes, we have a baseline model that uses crossed random intercepts to account for the multilevel structure of the experiment. The first random intercept is for participants, accounting for clustering due to repeated measures. This random intercept is then crossed with a second random intercept for response chains, which we nest inside of {[condition], item} combinations. In R syntax, the random effect structure was ... + (1|ParticipantID) +
(1|ItemCondition/ResponseChainID). Models were fit in the lme4 R package. We computed profile likelihood confidence intervals for coefficients using the confint R package. We used estimated marginal means (emmeans R package) to conduct model-adjusted F-tests, linear contrasts, predictions, and pairwise comparisons. We apply Holm-Bonferroni adjustments to pairwise comparison p-values. Our baseline ‘local’ model is:
$\text{variable}_{ijk}$ $=\beta_{0}+\beta_{1}\text{condition}_{j}+\beta_{2}\text{CreativityHuman}_{i}+$
$\quad\beta_{3}\text{AiRelCreate}_{i}+\beta_{4}\text{AiFeeling}_{i}+$
$\quad\beta_{5}\text{InterestGroup}_{i}+\beta_{6}\text{ConditionOrder}_{ijk}+$
$\quad\beta_{7}\text{LogDuration}_{ijk}+\beta_{8}\text{nSeedsPresent}_{ijk}+$
(1) $\quad\beta_{9}\text{TrialNo}_{jk}+u_{0i}+v_{0jk}+e_{ijk}$
@@
  • AiRelCreate is constructed as (self-perceived creativity to humans) - (self-perceived creativity to AI). Note that this is an implicit measure of AI’s (perceived) creativity relative to humans. For example, if you say you are more creative than 40% of humans and 60% of AI, then AiRelCreate = -20, as the implicit belief is that AI is less creative (-20 percentile points) than humans. Conversely, if you say you are more creative than 50% of humans but 30% of AI then the implicit belief is that AI is more creative (50% - 30% = +20) than humans.
@@
We balanced interest in testing experimental hypotheses that conditions differed by subgroups with caution around model overfitting. We considered interactions between the treatment condition and four potential moderators: self-perceived human creativity ($CreativityHuman$), AI - Human creativity ($AiRelCreate$), feeling towards AI ($AiFeeling)$, and interest group ($InterestGroup$). We first conducted likelihood ratio tests to test if adding each moderator improved our baseline model. Moderators were kept only if they significantly improved the fit $(p<0.05)$. See SI 7 for retained moderators. Then, we used emmeans to probe and interpret moderating effects.
B.2. Global